# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v2

@@ Abstract
Questionnaires in the behavioral and organizational sciences tend to be lengthy: survey measures comprising hundreds of items are the norm rather than the exception. However, literature suggests that the longer a questionnaire takes, the higher the probability that participants lose interest and start responding carelessly. Consequently, in long surveys a large number of participants may engage in careless responding, posing a major threat to internal validity. We propose a novel method for identifying the onset of careless responding (or an absence thereof) for each participant. It is based on combined measurements of multiple dimensions in which carelessness may manifest, such as inconsistency and invariability. Since a structural break in either dimension is potentially indicative of carelessness, the proposed method searches for evidence for changepoints along the combined measurements. It is highly flexible, based on machine learning, and provides statistical guarantees on its performance. An empirical application on data from a seminal study on the incidence of careless responding reveals that the reported incidence has likely been substantially underestimated due to the presence of respondents that were careless for only parts of the questionnaire. In simulation experiments, we find that the proposed method achieves high reliability in correctly identifying carelessness onset, discriminates well between careless and attentive respondents, and captures a variety of careless response types, even when a large number of careless respondents are present. Furthermore, we provide freely available open source software to enhance accessibility and facilitate adoption by empirical researchers.
@@ Introduction
According to the definition of Ward & Meade, (2023), “careless responding occurs when participants are not basing their response on the item content, and it can occur when a respondent does not read an item, does not understand an item, or is unmotivated to think about what the item is asking”.11endnote: 1Perhaps more commonly used, Huang et al., (2012) define careless responding as “a response set in which the respondent answers a survey measure with low or little motivation to comply with survey instructions, correctly interpret item content, and provide accurate responses”. We prefer the definition of Ward & Meade, (2023), as it highlights that careless responding may not be intentional (cf. Huang et al.,, 2012; Ward & Pond,, 2015; Ward et al.,, 2017). Also note that Huang et al., (2012) refer to careless responding as insufficient effort responding in their definition. Other synonyms for careless responding are participant inattention (Maniaci & Rogge,, 2014), inconsistent responding (Greene,, 1978), protocol invalidity (Johnson,, 2005), and random responding (e.g., Beach,, 1989; Berry et al.,, 1992; Emons,, 2008). Notably, Schroeders et al., (2022) point out that the latter—random responding—may be a misnomer since careless responding can also arise in non-random patterns (e.g., a recurring sequence of 1-2-3-4-5). Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as lower scale reliability, errors in statistical hypothesis testing, and deteriorated model fit (e.g., Arias et al.,, 2020; Huang et al.,, 2012, 2015; Kam & Meyer,, 2015). It is widely prevalent (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2015; Curran,, 2016; Meade & Craig,, 2012), suspected to be present in all questionnaire data (Ward & Meade,, 2023), and already incidence rates as low as 5% can be problematic (Arias et al.,, 2020; Credé,, 2010). Often viewed as merely an insidious source of measurement error, careless responding can be a valuable source of information about study participants and the study itself. For instance, carelessness has been linked to participant personality (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2020) and design factors such as study length (e.g., Bowling et al.,, 2021) and item wording. In particular negatively-worded items are prone to misresponses (e.g., Swain et al.,, 2008; Weijters & Baumgartner,, 2012) triggered by carelessness. Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al.,, 2015; Huang et al.,, 2015).
@@
Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance consistency indicators such as psychometric synonyms (Meade & Craig,, 2012), longstring indices (Johnson,, 2005), or multivariate outlier analyses (e.g., Curran,, 2016). We refer to Arthur et al., (2021), DeSimone et al., (2015), and Ward & Meade, (2023) for detailed reviews.22endnote: 2A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al.,, 2021). Such detection items are based on the presumption that an attentive respondent will respond in a specific way, while careless respondent may fail to do so. For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade & Craig,, 2012). A careless respondent may accidentally “agree” to this item as a consequence of inattention. Alternative types of detection items are instructed items and self-report items (Arthur et al.,, 2021). We discuss detection items and preventive measures against carelessness in detail in a later section. More recently, Schroeders et al., (2022), Alfons & Welz, (2024), and Ozaki, (2024) have studied machine learning techniques for identifying careless respondents. Still, commonly used screeners tend to suffer from three major limitations, which we discuss in the following.
First, extant methods commonly follow the implicit assumption that careless respondents respond carelessly to all questionnaire items or at least a vast majority of items. In particular, they indicate whether a respondent engaged in careless responding, but they typically do not provide any indication regarding which items are affected. Although there exist many comparative studies (e.g., Curran,, 2016; DeSimone & Harms,, 2018; Goldammer et al.,, 2020; Huang et al.,, 2012) on such screeners, their behavior under partial careless responding is not well-understood. This is at odds with the substantial evidence that many careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding due to fatigue or boredom as a survey progresses (Bowling et al.,, 2021; Brower,, 2018; Galesic & Bosnjak,, 2009; Gibson & Bowling,, 2020; Ward et al.,, 2017). In general, Ward & Meade, (2023) believe it is rare that careless respondents respond carelessly from the beginning to the end of a questionnaire.
Second, common detection methods are typically designed to detect a single specific type of carelessness. For instance, the longstring index of Johnson, (2005) counts the maximum number of consecutive identical responses, which is intended to capture straightlining behavior. However, careless responding may manifest in multiple distinctive ways (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023), implying that a screener that is designed to detect one type of carelessness may fail to detect other types. It is therefore recommended to combine multiple screeners to detect various types of careless respondents (e.g., Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
Third, cutoff values for flagging careless respondents are often ad-hoc and lack statistical foundation. Screeners typically calculate a score for each participant, with a participant being flagged as careless if their score is sufficiently extreme (i.e., beyond a certain cutoff value). Ideally, such a cutoff value results in high sensitivity (flagging true careless respondents as careless) and high specificity (not flagging true attentive respondents as careless). For certain screeners, various cutoff values have been suggested in the literature, but individual suggestions vary substantially. For instance, for the same screeners, Goldammer et al., (2020) argues for different cutoff values than Curran, (2016) or Johnson, (2005), which suggests that the results for specific cutoff values may depend on factors such as the population or the constructs being measured (cf. Meade & Craig,, 2012, p. 452). For a specific screener, it would be desirable to obtain a cutoff value with statistical guarantees on sensitivity and specificity that does not depend on population or study characteristics.
@@
In wake of these three issues with extant methods, we propose a novel screener for the identification of careless responding that is designed to overcome these issues. We achieve this by focusing on a more nuanced question. Motivated by the prominence of partial careless responding due to survey fatigue or boredom, we focus on identifying at which point in the survey a given participant becomes careless (if at all). Specifically, our method estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof. Hence, we call our method careless onset detection in extensive rating-scale surveys (CODERS). To capture various types of carelessness, CODERS can combine multiple dimensions of evidence for careless responding to construct a score that, for each item, measures if a given respondent has started responding carelessly by that item. It achieves this by searching for changepoints along a multidimensional series of indices that are potentially indicative of carelessness, such as measures of response consistency or variability.
@@
We argue that the notion of a changepoint is intuitive when studying partial carelessness: Once a participant starts responding carelessly, we expect a structural break in their responses. In particular, such a respondent may abandon content-based responding and resort to careless response types, while no such break occurs in the absence of carelessness. CODERS is highly flexible as it does not assume a statistical model, nor does it require predefining what types of careless response types may exist, and it is primarily intended for lengthy multi-scale surveys. Moreover, CODERS builds on nonparametric estimation theory for cutoff values and statistical guarantees regarding sensitivity and specificity.
Re-examining data from a a seminal study on careless responding in Big 5 measurements (Johnson,, 2005), we demonstrate novel insights that can be obtained by estimating onset items of partially careless responding. Most notably, we find strong evidence that Johnson, (2005) considerably underestimates the incidence of careless respondents due to hitherto undetected presence of partially careless respondents. Through extensive simulation experiments, we find that the proposed method CODERS accurately estimates carelessness onset (or a lack thereof) and thereby achieves excellent sensitivity and specificity in detecting careless responding.
To the best of our knowledge, CODERS is the first attempt to systematically detect the onset of careless responding. Assuming that participants respond accurately in the beginning of the survey, CODERS is designed to segment the responses of each respondent into a segment of accurate responses and—if a changepoint was identified—a segment of careless responses. With this knowledge, researchers can restrict their primary analyses to the segments of accurate responses (by treating the careless responses as missing values) without having to discard all responses of a partially careless respondent. In addition, researchers can separately study the segments of careless responses to build theory on and obtain better understanding of the nature of (partially) careless responding (cf. Gibbert et al.,, 2021, for theory-building based on outlying data points). Finally, we provide the freely available software package carelessonset (Welz & Alfons,, 2024), which implements CODERS in the statistical programming environment R (R Core Team,, 2024) can be downloaded from https://github.com/mwelz/carelessonset. As such, it is a useful and accessible tool for any researcher who is concerned with survey fatigue or careless responding in general.
@@
The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023). We expect partial careless responding to result in a change in respondent behavior in at least one of these dimensions. CODERS therefore offers a general framework for searching for a joint changepoint in any combination of measurements of these dimensions of carelessness. While the majority of this paper is focused on inconsistency and invariability, we revisit response times in a later section.
We propose to measure the first dimension, inconsistency, by means of reconstructions of observed responses (cf. Alfons & Welz,, 2024). These reconstructions are generated by an auto-associative neural network (henceforth autoencoder) (Kramer,, 1992), which learns response patterns that characterize attentive responding. We expect that random-like, content-independent responses cannot be learned well and are thus poorly reconstructed by the autoencoder, leading to a changepoint in reconstruction performance.
For measuring the second dimension, invariability, we propose a novel algorithm that is inspired by the famous longstring index of Johnson, (2005). Since long sequences of identical responses or constant response patterns are not expected in surveys that use positively and negatively coded items in a randomized order, we expect a changepoint in response variability once a respondent commences to respond carelessly through straightlining or pattern responding behavior.
Overall, CODERS attempts to capture evidence for the onset of carelessness by combining multiple indicators that are potentially indicative of such an onset, where the different indicators may capture different manifestations of carelessness. Conversely, if a respondent never becomes careless, we do not expect a changepoint in any dimension. Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
We provide a detailed description of our assumptions in Appendix A and a technical definition of CODERS in Appendix B.
@@
To achieve the goal of learning the internal structures of a dataset, it is crucial to avoid that an autoencoder simply passes the input data through its network architecture. Consequently, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf. Kramer,, 1992, and Chapter 14.6 in Goodfellow et al.,, 2016). The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest. In this context, the autoencoder aims to learn the patterns between the latent constructs in the low-dimensional space. For attentive responses, randomness that can be attributed to taking multiple measurements of the latent constructs is thereby filtered out. Since the amount of randomness from these multiple measurements should be relatively low for a well-designed questionnaire, attentive responses are expected to be reconstructed well from the learned patterns in the low-dimensional space of the latent constructs. Inconsistent careless responses, on the other hand, should contain little to no structure also in this low-dimensional space, and are therefore expected to be reconstructed poorly.
A schematic example of the network architecture of an autoencoder is given in Figure 1. Concretely, an autoencoder is a fully connected network whose nodes are organized in multiple layers. Each node takes a linear combination of the nodes from the previous layer and subsequently applies a so-called activation function in order to allow for nonlinear relationships. A crucial part of an autoencoder is the central layer, known as the bottleneck layer, which stores the low-dimensional representations of the input data. The network architecture of an autoencoder is symmetric around this bottleneck layer such that the can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again. It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer,, 1991).
@@
Note. The network consists of five fully connected layers and is symmetric around the central layer. Each dimension is represented by one node, and nodes within the same layer are aligned vertically. In each node, a so-called activation function is applied to a linear combination of the nodes from the previous layer. The first layer (input layer) represents the input variables (here $\bm{x}_{1},\dots,\bm{x}_{4}$). The second layer (mapping layer) prepares the data to be compressed in the third and central layer (bottleneck layer), which holds low-dimensional representations of the input data (here $\bm{\widehat{z}}_{1}$ and $\bm{\widehat{z}}_{2}$). The fourth layer (demapping layer) prepares the low-dimensional respresentations for reconstructing the input data in the final layer (output layer), yielding reconstructions in the original dimension (here $\bm{\widehat{x}}_{1},\dots,\bm{\widehat{x}}_{4}$).
@@
A thorough mathematical definition of autoencoders can be found in Appendix B.1. Although there are other design choices besides the number of nodes in the bottleneck layer (see Appendix B.1.2), our software provides default values that we recommend researchers to use. The only input required by the user is the number of constructs measured in the survey (for the number of nodes in the bottleneck layer).
@@
where $L_{j}$ denotes the number of answer categories of item $j$. Recall that we expect the reconstruction errors $\textsf{RE}_{ij}$ to be low in case of attentive responding but high in the presence of inconsistent careless responding. Hence, if participant $i$ starts providing inconsistent careless responses at item $k\in\{1,\dots,p\}$, we expect a changepoint at position $k$ in the participant’s series of $p$ reconstruction errors, as reconstruction errors are expected to be higher from item $k$ onward. An example can be found in the top panel of Figure 2(a).
Figure 2: Illustrative example output from CODERS using responses of a participant to a 300-item personality inventory.
(a) Two dimensions indicative of carelessness: autoencoder reconstruction errors (RE) in the top panel, and the LongstringPattern (LSP) sequence in the bottom panel.
(b) The per-response statistic of the changepoint detection test. The dashed horizontal line corresponds to the associated critical value at an extremely conservative significance level of 0.1%.
Note. The solid vertical lines indicate the identified carelessness onset, which is at the 148th item. Reconstruction errors are in general much larger after this item than before, indicating more random-like responses. The response data are taken from Johnson, (2005) and are discussed in detail in a later section.
@@
Define $l\textsf{-pattern}_{ij}$ as the number of consecutive items in the longest sequence of a recurring pattern of length $l$ in which the response of participant $i$ to item $j$ takes part, or assign the value 1 if the response is not part of a recurring pattern of length $l$. Consider the following two illustrative examples. First, for $l=2$, the response sequence $1-2-1-2-1-4-3-5-4-5-4$ will be assigned the 2-pattern sequence $4-4-4-4-1-1-1-4-4-4-4$ because the first four responses and last four responses comprise two occurrences of the distinct patterns $1-2$ and $5-4$, respectively, which are both of length $l=2$. The central three responses ($1-4-3$) are not part of any recurring pattern of length $l=2$, hence they are each assigned the value 1. Second, for a recurring pattern of length $l=1$ (i.e., consecutive identical responses), the response sequence $3-2-3-3-1-4-1-1-1$ is assigned the 1-pattern sequence $1-1-2-2-1-1-3-3-3$, since the subsequences of consecutive identical responses $3-3$ and $1-1-$ 1 - 1 comprise two and three responses, respectively.33endnote: 3The 1-pattern sequence can be seen as a generalization of the the longstring index of Johnson, (2005), which we recover by calculating $\max_{1\leq j\leq p}1\textsf{-pattern}_{ij}$ for each respondent $i=1,\dots,n$. In short, higher values in the $l$-pattern sequence capture a tendency to respond in fixed patterns of length $l$. If a participant starts engaging in invariable careless responding, we expect a changepoint in the $l$-pattern sequence (for some $l$) from relatively low to relatively high values, assuming that the order of items is randomized.
@@
for participant $i=1,\dots,n$ and item $j=1,\dots,p$. In this paper, we vary $l$ from 1 to the number of answer categories in the survey. The rationale behind this choice is that we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
We provide a detailed description of the LSP algorithm in Appendix B.2. An example of an LSP sequence is shown in the bottom panel of Figure 2(a).

### v3

@@ Abstract
Questionnaires in the behavioral and organizational sciences tend to be lengthy. However, literature suggests that survey length is a contributing factor to careless responding, with longer questionnaires yielding higher probability that participants start responding carelessly. Consequently, in long surveys a large number of participants may engage in careless responding, posing a major threat to internal validity. We propose a novel method for identifying the onset of careless responding (or an absence thereof) that searches for a changepoint in combined measurements of multiple dimensions in which carelessness may manifest, such as inconsistency and invariability. It is highly flexible, based on machine learning, and provides statistical guarantees for controlling the false positive rate. In simulation experiments, the proposed method achieves high accuracy in identifying carelessness onset and discriminates well between attentive and various types of careless responding, even when a large number of careless respondents are present. An empirical application highlights how identifying partial carelessness uncovers novel insights on careless responding behavior. Furthermore, we provide the freely available open source software package carelessonset to facilitate adoption by empirical researchers.
@@ Introduction
According to the definition of Ward & Meade, (2023), “careless responding occurs when participants are not basing their response on the item content, and it can occur when a respondent does not read an item, does not understand an item, or is unmotivated to think about what the item is asking”.11endnote: 1Perhaps more commonly used, Huang et al., (2012) define careless responding as “a response set in which the respondent answers a survey measure with low or little motivation to comply with survey instructions, correctly interpret item content, and provide accurate responses”. We prefer the definition of Ward & Meade, (2023), as it highlights that careless responding may not be intentional (cf. Huang et al.,, 2012; Ward & Pond,, 2015; Ward et al.,, 2017). Also note that Huang et al., (2012) refer to careless responding as insufficient effort responding in their definition. Other synonyms for careless responding are participant inattention (Maniaci & Rogge,, 2014), inconsistent responding (Greene,, 1978), protocol invalidity (Johnson,, 2005), and random responding (e.g., Beach,, 1989; Berry et al.,, 1992; Emons,, 2008). Notably, Schroeders et al., (2022) point out that the latter—random responding—may be a misnomer since careless responding can also arise in non-random patterns (e.g., a recurring sequence of 1-2-3-4-5). Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as lower scale reliability, errors in statistical hypothesis testing, and deteriorated model fit (e.g., Arias et al.,, 2020; Huang et al.,, 2012, 2015; Kam & Meyer,, 2015). It is widely prevalent (e.g., Bowling et al.,, 2016; Curran,, 2016; DeSimone et al.,, 2015; Meade & Craig,, 2012), suspected to be present in all questionnaire data (Ward & Meade,, 2023), and already incidence rates as low as 5% can be problematic (Arias et al.,, 2020; Credé,, 2010; Welz et al., 2024a, ; Welz et al., 2024b, ). Often viewed as merely an insidious source of measurement error, careless responding can be a valuable source of information about study participants and the study itself. For instance, carelessness has been linked to participant personality (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2020) and design factors such as study length (e.g., Bowling et al.,, 2021) and item wording. In particular negatively-worded items are prone to misresponses (e.g., Swain et al.,, 2008; Weijters & Baumgartner,, 2012) triggered by carelessness. Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al.,, 2015; Huang et al.,, 2015).
@@
Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance inconsistency indicators such as psychometric synonyms or antonyms (Meade & Craig,, 2012), invariability indicators such as the longstring index (Johnson,, 2005), or multivariate outlier analyses (e.g., Curran,, 2016). We refer to Arthur et al., (2021), DeSimone et al., (2015), and Ward & Meade, (2023) for detailed reviews.22endnote: 2A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al.,, 2021). Such detection items are based on the presumption that an attentive respondent will respond in a specific way, while careless respondent may fail to do so. For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade & Craig,, 2012). Alternative types of detection items are instructed items and self-report items (Arthur et al.,, 2021). More recently, Alfons & Welz, (2024), Ozaki, (2024), and Schroeders et al., (2022) have studied machine learning techniques for identifying careless respondents. Existing methods and their implementations in freely available software packages such as careless (Yentes & Wilhelm,, 2023) make screening for careless respondents a simple task for empirical researchers, thus elevating reliability and validity of research findings and playing an instrumental role in gaining insights into the phenomenon of careless responding. In addition, many popular screeners are intuitive and easy to understand. It should be noted that common indicators typically compute a single score for each participant, with respondents deemed careless and removed from subsequent analysis if their score exceeds a certain cutoff value. Although there exist many comparative studies on such screeners (e.g., Curran,, 2016; DeSimone & Harms,, 2018; Goldammer et al.,, 2020; Huang et al.,, 2012), their behavior under partially careless responding is unfortunately not well-understood. Furthermore, excluding partially careless respondents—particularly in lengthy surveys—may leave a lot of valuable information unused, namely the respective parts of the survey in which they responded attentively.
We therefore aim to address a more nuanced question: Can we identify at which point in the survey participants start responding carelessly, if at all? Indeed, there is evidence that many—though not all—careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding as the survey progresses, for instance due to fatigue or boredom (Bowling et al.,, 2021; Brower,, 2018; Galesic & Bosnjak,, 2009; Gibson & Bowling,, 2020; Ward et al.,, 2017); cf. also the literature on back random responding (e.g., Clark et al.,, 2003) or partial random responding (e.g. Berry et al.,, 1992; Gallen & Berry,, 1997). Successful identification of the onset of careless responding can therefore yield valuable, complementary insights to those from existing screeners. For subsequent analysis, the issue of partially careless responding may be transformed into a missing data problem. Rather than excluding partially careless respondents altogether, researchers can restrict their primary analyses to the parts of attentive responses, treating the careless responses as missing values. Common guidelines for handling missing data (e.g., Newman,, 2014) can thereby be followed. In addition, accurately identifying the onset of careless responding may uncover implications for survey design (e.g., if many participants start responding carelessly at a similar point in the survey, it may be beneficial to rephrase or replace those items, or to place an intervention) and may lead to new theory and a deeper understanding of careless responding behavior (e.g., potentially linking certain personality traits to a tendency for an earlier or later onset of careless responding; cf. Gibbert et al.,, 2021, for theory-building based on outlying data).
@@
For the purpose of identifying at which point in the survey a given participant becomes careless, if at all, we introduce a novel method called careless onset detection in extensive rating-scale surveys (CODERS). Specifically, CODERS estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof. An important issue is that careless responding may manifest in multiple distinctive ways (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023). To capture various types of carelessness, CODERS can combine multiple dimensions of evidence for careless responding to construct a score that, for each item, measures if a given respondent has started responding carelessly by that item. It achieves this by searching for changepoints along a multidimensional series of indices that are potentially indicative of carelessness, such as measures of response inconsistency or invariability.
@@
We argue that the notion of a changepoint is intuitive when studying partial carelessness: if a participant starts responding carelessly, we expect a structural break in their responses. In particular, such a respondent may abandon content-based responding and resort to careless response types from the changepoint onward. CODERS is highly flexible as it does not assume a statistical model, nor does it require predefining what types of careless response types may exist, and it is primarily intended for lengthy multi-scale surveys. Moreover, CODERS builds on nonparametric estimation theory for cutoff values with statistical guarantees regarding the false positive rate (i.e., the probability of attentive respondents being incorrectly flagged as partially careless).
To the best of our knowledge, CODERS is the first attempt to systematically detect the onset of careless responding. Through extensive simulation experiments, we find that CODERS accurately estimates carelessness onset (or a lack thereof) and achieves excellent sensitivity and specificity in detecting partially careless responding. Crucially, CODERS should not be viewed as a replacement of existing screeners for careless respondents, but as a complementary tool for a related but more nuanced task, as highlighted by an empirical application. Re-examining data from a a seminal study on careless responding in Big Five personality measurements (Johnson,, 2005), we demonstrate novel insights that can be obtained by estimating onset items of partially careless responding. Reflecting on those insights, we subsequently discuss limitations and potential extensions of CODERS. Our concluding discussion highlights how CODERS complements various existing approaches for the detection of careless respondents. Finally, we provide the freely available software package carelessonset (Welz & Alfons,, 2024), which implements CODERS in the statistical programming environment R (R Core Team,, 2024) and can be downloaded from https://github.com/mwelz/carelessonset. As such, it is a useful and accessible tool for any researcher who is concerned with survey fatigue or careless responding in general.
@@
The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023). We expect partially careless responding to result in a change in respondent behavior in at least one of these dimensions. CODERS therefore offers a general framework for searching for a joint changepoint in any combination of measurements of these dimensions of carelessness. While the majority of this paper is focused on inconsistency and invariability, we revisit response times in a later section. As used here, CODERS thus consists of three main components:
  1. 1.
    For each participant, a nonparametric changepoint detection test (Shao & Zhang,, 2010) is applied to a combined series of participant-item level measurements of careless responding.
  2. 2.
    As a measurement of inconsistency, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer,, 1992) are used. An autoencoder can be viewed as a nonlinear generalization of principal component analysis (PCA), and the (participant-item level) reconstruction errors can be interpreted as the difference between observed responses and what is expected under attentive responding.
  3. 3.
    As a measurement of invariability, we propose a novel algorithm called LongStringPattern (LSP). LSP can be viewed as a generalization of the famous longstring index of Johnson, (2005) to the participant-item level, also allowing for recurring patterns involving multiple answer categories.
Regarding the first dimension, inconsistency, the autoencoder learns response patterns that characterize attentive responding (cf. Alfons & Welz,, 2024). We expect that random-like, content-independent responses cannot be learned well and are thus poorly reconstructed by the autoencoder, leading to a changepoint in reconstruction performance. Regarding the second dimension, invariability, long sequences of identical responses or fixed response patterns are not expected in well-designed surveys (e.g., using positively and negatively coded items, randomized in some form). Hence, we expect a changepoint in response variability once a respondent commences to respond carelessly through straightlining or pattern responding behavior. To obtain a granular estimate of carelessness onset, both reconstruction errors and LSP sequences are measured on the participant-item level so that a changepoint can be detected on the item-level for each participant.
Overall, CODERS attempts to capture evidence for the onset of carelessness by combining multiple indicators that are potentially indicative of such an onset, where the different indicators may capture different manifestations of carelessness. Conversely, if a respondent never becomes careless, we do not expect a changepoint in any dimension. Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023). We provide a detailed description of our assumptions in Appendix A and additional details on CODERS in Appendix B.
@@
To achieve the goal of learning the internal structures of a dataset, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf. Kramer,, 1992, and Chapter 14.6 in Goodfellow et al.,, 2016). The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest, such as attitudes, sentiments, or personality traits. In this context, the autoencoder aims to learn the patterns between the latent constructs in the low-dimensional space. For attentive responses, randomness that can be attributed to taking multiple measurements of the latent constructs is thereby filtered out. Since the amount of randomness from these multiple measurements should be relatively low for a well-designed questionnaire, attentive responses are expected to be reconstructed well from the learned patterns in the low-dimensional space of the latent constructs. Inconsistent careless responses, on the other hand, should contain little to no structure also in this low-dimensional space, and are therefore expected to be reconstructed poorly.
@@
Note. The network consists of five fully connected layers and is symmetric around the central layer. Each dimension is represented by one node, and nodes within the same layer are aligned vertically. In each node, a so-called activation function is applied to a linear combination of the nodes from the previous layer. The first layer (input layer) represents the input variables (here $\bm{x}_{1},\dots,\bm{x}_{4}$). The second layer (mapping layer) prepares the data to be compressed in the third and central layer (bottleneck layer), which holds low-dimensional representations of the input data (here $\bm{\widehat{z}}_{1}$ and $\bm{\widehat{z}}_{2}$). The fourth layer (demapping layer) prepares the low-dimensional representations for reconstructing the input data in the final layer (output layer), yielding reconstructions in the original dimension (here $\bm{\widehat{x}}_{1},\dots,\bm{\widehat{x}}_{4}$).
A schematic example of the network architecture of an autoencoder is given in Figure 1. Concretely, an autoencoder is a fully connected network whose nodes are organized in multiple layers. Each node takes a linear combination of the nodes from the previous layer and subsequently applies a so-called activation function in order to allow for nonlinear relationships. A crucial part of an autoencoder is the central layer, known as the bottleneck layer, which stores the low-dimensional representations of the input data. The network architecture of an autoencoder is symmetric around this bottleneck layer such that it can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again. It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer,, 1991).
@@
Further intuition and a thorough mathematical definition of autoencoders can be found in Appendix B.2. Although there are other design choices besides the number of nodes in the bottleneck layer (see Appendix B.2.3), our software implementation in package carelessonset (Welz & Alfons,, 2024) provides default values that we recommend researchers to use. The only input required by the user is the number of constructs measured in the survey (for the number of nodes in the bottleneck layer).
@@
where $L_{j}$ denotes the number of answer categories of item $j$. Recall that we expect the reconstruction errors $\textsf{RE}_{ij}$ to be low in case of attentive responding but high in the presence of inconsistent careless responding. Hence, if participant $i$ starts providing inconsistent careless responses at item $k\in\{1,\dots,p\}$, we expect a changepoint at position $k$ in the participant’s series of $p$ reconstruction errors, as reconstruction errors are expected to be higher from item $k$ onward. An example can be found in Figure 2, where we show simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts choosing answer categories completely at random from the 172nd item, resulting in systemically larger autoencoder reconstruction errors from that item onward.
Figure 2: Simulated illustrative example of autoencoder reconstruction errors of an inconsistent partially careless respondent to a 300-item questionnaire comprising five-point Likert scales.
Note. The simulated participant started choosing answer categories completely at random from the 172nd item onward.
@@
Define $l\textsf{-pattern}_{ij}$ as the number of consecutive items in the longest sequence of a recurring pattern of length $l$ in which the response of participant $i$ to item $j$ takes part, or assign the value 1 if the response is not part of a recurring pattern of length $l$. Consider the following two illustrative examples. First, for $l=2$, the response sequence $1-2-1-2-1-4-3-5-4-5-4$ will be assigned the 2-pattern sequence $4-4-4-4-1-1-1-4-4-4-4$ because the first four responses and last four responses comprise two occurrences of the distinct patterns $1-2$ and $5-4$, respectively, which are both of length $l=2$. The central three responses ($1-4-3$) are not part of any recurring pattern of length $l=2$, hence they are each assigned the value 1. Second, for a recurring pattern of length $l=1$ (i.e., consecutive identical responses), the response sequence $3-2-3-3-1-4-1-1-1$ is assigned the 1-pattern sequence $1-1-2-2-1-1-3-3-3$, since the subsequences of consecutive identical responses $3-3$ and $1-1-$ 1 - 1 comprise two and three responses, respectively.33endnote: 3The 1-pattern sequence can be seen as a generalization of the the longstring index of Johnson, (2005), which we recover by calculating $\max_{1\leq j\leq p}1\textsf{-pattern}_{ij}$ for each respondent $i=1,\dots,n$. In short, higher values in the $l$-pattern sequence capture a tendency to respond in fixed patterns of length $l$. If a participant starts engaging in invariable careless responding, we expect a changepoint in the $l$-pattern sequence (for some $l$) from relatively low to relatively high values, assuming that the survey design precludes such patterns in attentive responses.
@@
for participant $i=1,\dots,n$ and item $j=1,\dots,p$. We call the series of within-participant values $\textsf{LSP}_{ij}$ across all items a LongStringPattern (LSP) sequence. In this paper, we vary $l$ from 1 to the number of answer categories in the survey. The rationale behind this choice is that a careless respondent may repeat, e.g., the pattern $1-2-3-4-5$ in case of five answer categories. At the same time, we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
We provide a detailed description of the LSP algorithm in Appendix B.3. An example of an LSP sequence is shown in Figure 3, which shows simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts responding carelessly using the recurring pattern $3-4$ from the 260th item, resulting in systematically larger LSP values from that item onward.
Figure 3: Simulated illustrative example of a LongStringPattern sequence of an invariable partially careless respondent to a 300-item questionnaire comprising five-point Likert scales.
Note. The simulated participant started responding carelessly with the recurring pattern $3-4$ from the 260th item onward.